# Approximation schemes for covering and packing mixed-integer programs with a fixed number of constraints

### v1

@@ 2 Decomposition into multidimensional knapsack problems
where (LABEL:eq:w-definition) follows from the definition of $w$, and (14) holds because, for each $j\in[m]$ and $i\in[n]\setminus\{g(j)\}$, we have $x_{ij}=\ell_{ij}y_{i}$ if $i\in L_{g(j),j}$, $x_{ij}=c_{ij}y_{i}$ if $i\in C_{g(j),j}$, and $x_{ij}=\alpha^{g}_{j}+\ell_{g(j),j}$ otherwise. Finally, it is clear that any feasible solution of (6)–(12) induces a feasible solution of (2)–(4) with the same value. Therefore, (5)–(12) is a correct formulation of p. ∎
@@ 3 A polynomial-time approximation scheme
Next, we design a polynomial-time approximation scheme for p (with $m$ fixed), which is based on the decomposition presented in the previous section. friezeApproximationAlgorithmsMdimensional1984 propose a PTAS for the Multidimensional Knapsack Cover problem without continuous variables. In what follows, we extend their PTAS to include exactly one continuous variable per dimension. The analysis of the proposed algorithm follows the methodology of friezeApproximationAlgorithmsMdimensional1984.
@@
The proof is very similar to the proof in friezeApproximationAlgorithmsMdimensional1984. We give it here for completeness.
@@
The algorithm’s runtime is dominated by solving the linear program on line 10 for each of the $\mathcal{O}(\eta^{k})$ subsets $S$. cohenSolvingLinearPrograms2021 present the current state-of-the-art method to solve a linear program in $\mathcal{O}\left(M(\eta)\log\eta\log\left(\eta/2^{-\mathcal{O}(L)}\right)\right)$ time where $M(\eta)\sim\eta^{2.38}$ is the cost of matrix multiplication and inversion, and $L=\mathcal{O}\left(\log(\eta+\mathinner{\!\left\lVert\bar{d}\right\rVert}_{\infty}+\mathinner{\!\left\lVert\bar{f}\right\rVert}_{\infty}+\mathinner{\!\left\lVert\bar{v}\right\rVert}_{\infty})\right)$. A consequence of using their method is that while optimal primal and dual solutions are obtained, an optimal basis is not necessarily identified. Therefore, we apply the algorithm due to megiddoFindingPrimalDualOptimal1991 to compute an optimal basis in strongly polynomial time from an optimal primal-dual solution pair. ∎
@@
Proposition 2.4 shows a decomposition of p into $\mathcal{O}(n^{m})$ instances of the multidimensional knapsack cover problem with a single continuous variable per dimension. One can run Algorithm 1 on each of these instances and output the cheapest solution. By Corollary 2.5 and Theorem 3.1, this leads to a PTAS for p when $m$ is fixed. ∎
Theorem 3.2 gives, in a sense, the best polynomial-time approximation one can design to p since the multidimensional knapsack cover problem – even without continuous variables – does not admit any fully polynomial-time approximation scheme unless $\mathsf{P}=\mathsf{NP}$ (magazine1984note).
3.1 Extension to packing problems
@@ 3.2 Extension to assignment problems
where $m$ and $n$ are positive integers, $v,\ell,c,f\in\mathbb{Z}_{\geq}^{nm}$ and $d\in\mathbb{Z}_{\geq}^{m}$ denote the nonnegative integer parameters. It is assumed, without loss of generality, that $d_{j}\geq c_{ij}\geq\ell_{ij}$ for all $i\in[n]$ and $j\in[m]$. Note that q includes relevant problems as the Generalized Assignment Problem (gap). In particular, when $v_{ij}=0$ and $c_{ij}=\ell_{ij}$ for all $i\in[n]$ and $j\in[m]$q is equivalent to gap, which is known to be $\mathsf{APX}$-hard (chekuri2005polynomial). On the positive side, for any $\epsilon>0$, there is an approximation for gap with ratio $(e/(e-1)+\epsilon)\approx(1.582+\epsilon)$ using an LP-rounding approach, which is currently the best known approximation for this problem (fleischerTightApproximationAlgorithms2006).
@@
Furthermore, this approach also yields a PTAS for the covering version of q. Moreover, by replacing $\sum_{j\in[m]}y_{ij}\leq 1$ with $\sum_{j\in[m]}y_{ij}\leq t_{i}$ for all $i\in[n]$ for a parameter $t\in[n]^{n}$, we can design a PTAS for a family of related problems using the same core method. Observe that we obtain q when $t_{i}=1$ for every $i\in[n]$, and the packing version of p when $t_{i}=n$ for all $i\in[n]$.
@@ 4 FPTAS and approximate formulation for $m=1$
In this section, we design a fully polynomial-time algorithm for p when $m=1$, denoted as p1. This problem generalizes the knapsack cover problem as this occurs when $\ell_{i}=c_{i}$ for all $i\in[n]$. Hence p1 is clearly $\mathsf{NP}$-hard. Theorem 3.2 already proves that a PTAS exists for this problem. From the introduction we also know that vanhoeselFullyPolynomialApproximation2001 present an FPTAS for the single-item capacitated economic lot-sizing problem (which includes the knapsack cover problem) that is more general than p1 in its consideration of backlogging and holding costs. However, their model does not include lower bounds. This distinction in problem scope is accompanied by a methodological divergence as their approach does not rely on a decomposition into knapsack sets. The current state-of-the-art approximation algorithm for the knapsack problem is a randomized pseudo-polynomial algorithm by qizheng2024simple, which achieves a time complexity of $\mathcal{O}(n^{3/2}\times\min\{w_{\max},p_{\max}\})$. However, it is not obvious how the introduction of a continuous variable affects this result, a question that warrants its own investigation.
@@
The running-time of Algorithm 3 on input $I^{\prime}$ takes $\mathcal{O}(n^{2}f^{\prime}_{\max})$ time , with $f^{\prime}_{\max}=\max_{i\in[\eta]}f^{\prime}_{i}$. It follows from the scaling in lines 18-20 that $f^{\prime}_{\max}\leq f_{\max}/\lambda+1\leq\eta\operatorname{poly}(\eta)/\epsilon+1$. Therefore, Algorithm 3 runs in $\mathcal{O}\left(\left(\eta\operatorname{poly}(\eta)/\epsilon+1\right)\eta^{2}\right)$ time. ∎
The hypothesis in Theorem 4.1 on the polynomial bound on the ratio $f_{\max}/f_{\min}$ is due to fact that the optimal value of an instance of 1-mkc can be arbitrarily smaller than $f_{\max}$ in general. As an example, consider an instance of 1-mkc with $\eta=2$, $\bar{d}=1$, $f_{1}=2$, $w_{1}=1$, $f_{2}=100$, $w_{2}=100$, and $\bar{v}=100$. Here, $f_{\max}=100$ while the optimal value equals $2$. A key direction for future research is to eliminate the assumption on the polynomial bound on the ratio $f_{\max}/f_{\min}$. We hypothesize that an approach similar to that of bienstockTighteningSimpleMixedinteger2012 could be adapted for this purpose, but formally establishing this remains an open challenge.
@@
Proposition 2.4 shows a decomposition of p1 into $\mathcal{O}(n)$ instances of the knapsack cover problem with a single continuous variable. One can run Algorithm 3 on each of these instances and output the cheapest solution. By Corollary 2.5 and Theorem 4.1, this leads to an FPTAS for p1 when for every resulting knapsack cover problem with a single continuous variable, $f_{\max}\geq f_{\min}\geq f_{\max}/\operatorname{poly}(n)$, where $f_{\max}=\max_{i\in[\eta]}\bar{f}_{i}$ and $f_{\min}=\min\left\{\min_{i\in[\eta]}\bar{f}_{i},\>\bar{v}\bar{d}\right\}$, and $\operatorname{poly}(\eta)\geq 1$ is any polynomial function on $\eta$. It is clear that this is guaranteed by the assumption on the input data of p. The total running time of this algorithm is $\mathcal{O}(n^{4}\operatorname{poly}(n)/\epsilon)$. ∎
@@
An important consequence of Theorem 3.2 and the equivalence of optimization and separation is that an $\epsilon$-approximate formulation must exist for p1. Since Proposition 2.4 decomposes p1 into $\mathcal{O}(n)$ instances of 1-mkc, we can directly leverage the results of bienstockTighteningSimpleMixedinteger2012, who obtain a provably tight, polynomially large formulation for the standard knapsack cover problem. We extend their methodology to incorporate a continuous variable, defining a linear programming formulation whose optimal value is at most $(1+\epsilon)$ times the optimal value of 1-mkc.
@@
Proposition 2.4 shows a decomposition of p1 into $\mathcal{O}(n)$ instances of the knapsack cover problem with a single continuous variable. For each of these problems, one may create an $\epsilon$-approximate formulation with $\mathcal{O}\left(\epsilon^{(\log\epsilon/\log(1+\epsilon))-1}n^{2}\right)$ variables and constraints, and combine them as in the formulation presented in Section 2. This leads to an $\epsilon$-approximate formulation for p1 with $\mathcal{O}\left(\epsilon^{(\log\epsilon/\log(1+\epsilon))-1}n^{3}\right)$ variables and constraints. ∎
The proof of the $\epsilon$-approximate formulation for the knapsack cover problem with a single continuous variable can be found in the appendix as the approach and structure adheres closely to what is shown by bienstockTighteningSimpleMixedinteger2012.
The existence of an $\epsilon$-approximate formulation for the packing version of p1 is guaranteed in principle by the equivalence of optimization and separation. However, such a formulation is not immediately available. A key complication arises from the presence of a continuous variable after the decomposition of p1. While bienstockApproximateFormulations012008 successfully derived an approximate formulation for the standard 0-1 knapsack set, it is unclear how to adapt the techniques to handle an additional continuous variable. Resolving this gap, and thus developing an explicit $\epsilon$-approximate formulation for the packing variant of p1, is a significant open challenge.
@@ 5 Perfect compact formulation for $m=1$ and uniform bounds
In this section, we present a compact and tight formulation for p when $m=1$ and the bounds are uniform, that is, $\ell_{i}=\ell$ and $c_{i}=c$ for all $i\in[n]$. This special case is denoted as pu. As mentioned in the literature review, pu can be solved in polynomial time (hellionPolynomialTimeAlgorithm2012a). However, no tight formulation of the problem is known. Hence the proposed perfect compact formulation contributes in this direction.
constantinoLowerBoundsLotSizing1998a shows that solving pu, where $\ell_{i}$ can vary with $i\in[n]$, is in general $\mathsf{NP}$-hard. This paper also analyzes a setting identical to pu and obtains two families of valid inequalities. Using both families, it is possible to describe the convex hull when $c\geq d$. agraPolyhedralDescriptionInteger2006a study the case where the $y$ variables are general integers and unbounded.
dashContinuousKnapsackSet2016a study the convex hull of the continuous knapsack set. Among other results, they prove that the convex hull with two integer and one bounded continuous variable is described using inequalities for a knapsack set with two integer variables and the ones for a continuous knapsack set with two integer and one continuous variable that is unbounded. We use a special case of this result below.
@@
By Theorem 4.4. in dashContinuousKnapsackSet2016a, we know that $\operatorname{conv}\{(\alpha,\gamma)\in\mathbb{R}_{\geq}\times\mathbb{Z}_{\geq}\mathrel{\mathop{\ordinarycolon}}\alpha+\gamma\geq\delta,\alpha\leq\sigma\}$ is the intersection of three sets, namely, $\operatorname{conv}\{(\alpha,\gamma)\in\mathbb{R}_{\geq}\times\mathbb{Z}_{\geq}\mathrel{\mathop{\ordinarycolon}}\alpha+\gamma\geq\delta\}$, $\{(\alpha,\gamma)\in\mathbb{R}\times\mathbb{R}\mathrel{\mathop{\ordinarycolon}}\alpha\leq\sigma\}$ and $\operatorname{conv}\{(\alpha,\gamma)\in\mathbb{R}\times\mathbb{Z}_{\geq}\mathrel{\mathop{\ordinarycolon}}\gamma\geq\delta-\sigma\}$. In the nontrivial case, the mixed integer rounding inequality $\alpha\geq(\delta-\lfloor\delta\rfloor)(\lceil\delta\rceil-\gamma)$ is needed to describe the first set, and the integer rounding inequality $\gamma\geq\lceil\delta-\sigma\rceil$ is needed to describe the third set (see, e.g., wolsey2020integer). Adding an upper bound of $\nu$ on the integer variable $\gamma$ does not create fractional extreme points.
Now we know that in every extreme point of the set defined by the trivial inequalities and the valid inequalities (26) and (25), $\sum_{i=1}^{\nu}\psi_{i}$ is integer. If $\psi$ is fractional, then it has at least two entries that are fractional and such a point cannot be an extreme point. ∎
@@
$x_{g}^{gb}+\sum_{i=g+1}^{n}\left(d-(b+1)\ell-\left\lfloor\frac{(d-(b+1)\ell}{c}\right\rfloor c\right)y^{gb}_{i}$
$\hskip 28.45274pt\geq\left(\ell+\left\lceil\frac{d-(b+1)\ell}{c}\right\rceil\times\left(d-(b+1)\ell-\left\lfloor\frac{(d-(b+1)\ell}{c}\right\rfloor c\right)\right)z^{gb}$ $\forall g\in\mathcal{G},b\in B^{g},$
@@
$X^{g1}(b)=\left\{(x,y)\in\mathbb{R}^{n}\times\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}\sum_{i=1}^{g-1}y_{i}=b,y_{g}=1\right\},$
@@
after setting $\nu=n-g$, $\delta=\frac{d-(b+1)\ell}{c}$, $\sigma=\frac{c-\ell}{c}$ and $\alpha=\frac{x_{g}-\ell}{c}$ in Lemma 5.1. Using the result on the convex hull of the union of polyhedra (conforti2008compact; balas1998disjunctive), we obtain the tight formulation. ∎

### v2

@@ 2 Decomposition into multidimensional knapsack problems
where (LABEL:eq:w-definition) follows from the definition of $w$, and (14) holds because, for each $j\in[m]$ and $i\in[n]\setminus\{g(j)\}$, we have $x_{ij}=\ell_{ij}y_{i}$ if $i\in L_{g(j),j}$, $x_{ij}=c_{ij}y_{i}$ if $i\in C_{g(j),j}$, and $x_{ij}=\alpha^{g}_{j}+\ell_{g(j),j}$ otherwise. Finally, it is clear that any feasible solution of (6)–(12) induces a feasible solution of (2)–(4) with the same value. Therefore, (5)–(12) is a correct formulation of p.
@@ 3 A polynomial-time approximation scheme
Next, we design a polynomial-time approximation scheme for p (with $m$ fixed), which is based on the decomposition presented in the previous section. Frieze and Clarke Frieze and Clarke (1984) propose a PTAS for the Multidimensional Knapsack Cover problem without continuous variables. In what follows, we extend their PTAS to include exactly one continuous variable per dimension. The analysis of the proposed algorithm follows the methodology of Frieze and Clarke (1984).
@@
The proof is very similar to the proof in Frieze and Clarke (1984). We give it here for completeness.
@@
The algorithm’s runtime is dominated by solving the linear program on line 10 for each of the $\mathcal{O}(\eta^{k})$ subsets $S$. Cohen et al. (2021) present the current state-of-the-art method to solve a linear program in $\mathcal{O}\left(M(\eta)\log\eta\log\left(\eta/2^{-\mathcal{O}(L)}\right)\right)$ time where $M(\eta)\sim\eta^{2.38}$ is the cost of matrix multiplication and inversion, and $L=\mathcal{O}\left(\log(\eta+\mathinner{\!\left\lVert\bar{d}\right\rVert}_{\infty}+\mathinner{\!\left\lVert\bar{f}\right\rVert}_{\infty}+\mathinner{\!\left\lVert\bar{v}\right\rVert}_{\infty})\right)$. A consequence of using their method is that while optimal primal and dual solutions are obtained, an optimal basis is not necessarily identified. Therefore, we apply the algorithm due to Megiddo Megiddo (1991) to compute an optimal basis in strongly polynomial time from an optimal primal-dual solution pair.
@@
Proposition 2.3 shows a decomposition of p into $\mathcal{O}(n^{m})$ instances of the multidimensional knapsack cover problem with a single continuous variable per dimension. One can run Algorithm 1 on each of these instances and output the cheapest solution. By Corollary 2.4 and Theorem 3.1, this leads to a PTAS for p when $m$ is fixed.
Theorem 3.2 gives, in a sense, the best polynomial-time approximation one can design to p since the multidimensional knapsack cover problem – even without continuous variables – does not admit any fully polynomial-time approximation scheme unless $\mathsf{P}=\mathsf{NP}$ (Magazine and Chern, 1984).
4 A polynomial-time approximation scheme to packing problems
@@
where parameters are defined as before.
@@
Furthermore, the same approach extends to the covering version of q, yielding a PTAS in the setting where (25) and (26) are $\geq$-constraints and the objective function is minimized. Moreover, by replacing $\sum_{j\in[m]}y_{ij}\leq 1$ with $\sum_{j\in[m]}y_{ij}\leq t_{i}$ for all $i\in[n]$ for a parameter $t\in[n]^{n}$, we obtain a more general version of q which also admits a PTAS using the same core method proposed in the previous sections. Observe that the packing version of p corresponds to the case where $t_{i}=n$ for all $i\in[n]$. Interestingly, the cover version of q is not a generalization of p.
@@ 5 FPTAS and approximate formulation for $m=1$
In this section, we design a fully polynomial-time algorithm for p when $m=1$, denoted as p1. This problem generalizes the knapsack cover problem as this occurs when $\ell_{i}=c_{i}$ for all $i\in[n]$. Hence p1 is clearly $\mathsf{NP}$-hard. Theorem 3.2 already proves that a PTAS exists for this problem. From the introduction we also know that Van Hoesel and Wagelmans Van Hoesel and Wagelmans (2001) present an FPTAS for the single-item capacitated economic lot-sizing problem (which includes the knapsack cover problem) that is more general than p1 in its consideration of backlogging and holding costs. However, their model does not include lower bounds. This distinction in problem scope is accompanied by a methodological divergence as their approach does not rely on a decomposition into knapsack sets. The current state-of-the-art approximation algorithm for the knapsack problem is a randomized pseudo-polynomial algorithm by He and Xu He and Xu (2024), which achieves a time complexity of $\mathcal{O}(n^{3/2}\times\min\{w_{\max},p_{\max}\})$ where $w_{\max}$ represents the maximum weight and $p_{\max}$ the maximum profit of an item. However, it is not obvious how the introduction of a continuous variable affects this result, a question that warrants its own investigation.
@@
The running-time of Algorithm 3 on input $I^{\prime}$ takes $\mathcal{O}(n^{2}f^{\prime}_{\max})$ time , with $f^{\prime}_{\max}=\max_{i\in[\eta]}f^{\prime}_{i}$. It follows from the scaling in lines 18-20 that $f^{\prime}_{\max}\leq f_{\max}/\lambda+1\leq\eta\operatorname{poly}(\eta)/\epsilon+1$. Therefore, Algorithm 3 runs in $\mathcal{O}\left(\left(\eta\operatorname{poly}(\eta)/\epsilon+1\right)\eta^{2}\right)$ time.
The hypothesis in Theorem 5.1 on the polynomial bound on the ratio $f_{\max}/f_{\min}$ is due to fact that the optimal value of an instance of 1-mkc can be arbitrarily smaller than $f_{\max}$ in general. As an example, consider an instance of 1-mkc with $\eta=2$, $\bar{d}=1$, $f_{1}=2$, $w_{1}=1$, $f_{2}=100$, $w_{2}=100$, and $\bar{v}=100$. Here, $f_{\max}=100$ while the optimal value equals $2$. A key direction for future research is to eliminate the assumption on the polynomial bound on the ratio $f_{\max}/f_{\min}$. We hypothesize that an approach similar to that of Bienstock and McClosky (2012) could be adapted for this purpose, but formally establishing this remains an open challenge.
@@
Proposition 2.3 shows a decomposition of p1 into $\mathcal{O}(n)$ instances of the knapsack cover problem with a single continuous variable. One can run Algorithm 3 on each of these instances and output the cheapest solution. By Corollary 2.4 and Theorem 5.1, this leads to an FPTAS for p1 when for every resulting knapsack cover problem with a single continuous variable, $f_{\max}\geq f_{\min}\geq f_{\max}/\operatorname{poly}(n)$, where $f_{\max}=\max_{i\in[\eta]}\bar{f}_{i}$ and $f_{\min}=\min\left\{\min_{i\in[\eta]}\bar{f}_{i},\>\bar{v}\bar{d}\right\}$, and $\operatorname{poly}(\eta)\geq 1$ is any polynomial function on $\eta$. It is clear that this is guaranteed by the assumption on the input data of p. The total running time of this algorithm is $\mathcal{O}(n^{4}\operatorname{poly}(n)/\epsilon)$.
@@
An important consequence of Theorem 3.2 and the equivalence of optimization and separation is that an $\epsilon$-approximate formulation must exist for p1. Since Proposition 2.3 decomposes p1 into $\mathcal{O}(n)$ instances of 1-mkc, we can directly leverage the results of Bienstock and McClosky (2012), who obtain a provably tight, polynomially large formulation for the standard knapsack cover problem. We extend their methodology to incorporate a continuous variable, defining a linear programming formulation whose optimal value is at most $(1+\epsilon)$ times the optimal value of 1-mkc.
@@
Proposition 2.3 shows a decomposition of p1 into $\mathcal{O}(n)$ instances of the knapsack cover problem with a single continuous variable. For each of these problems, one may create an $\epsilon$-approximate formulation with $\mathcal{O}\left(\epsilon^{(\log\epsilon/\log(1+\epsilon))-1}n^{2}\right)$ variables and constraints, and combine them as in the formulation presented in Section 2. This leads to an $\epsilon$-approximate formulation for p1 with $\mathcal{O}\left(\epsilon^{(\log\epsilon/\log(1+\epsilon))-1}n^{3}\right)$ variables and constraints.
The proof of the $\epsilon$-approximate formulation for the knapsack cover problem with a single continuous variable can be found in the appendix as the approach and structure adheres closely to what is shown by Bienstock and McClosky Bienstock and McClosky (2012).
The existence of an $\epsilon$-approximate formulation for the packing version of p1 is guaranteed in principle by the equivalence of optimization and separation. However, such a formulation is not immediately available. A key complication arises from the presence of a continuous variable after the decomposition of p1. While Bienstock Bienstock (2008) successfully derived an approximate formulation for the standard 0-1 knapsack set, it is unclear how to adapt the techniques to handle an additional continuous variable. Resolving this gap, and thus developing an explicit $\epsilon$-approximate formulation for the packing variant of p1, is a significant open challenge.
@@ 6 Perfect compact formulation for $m=1$ and uniform bounds
In this section, we present a compact and tight formulation for p when $m=1$ and the bounds are uniform, that is, $\ell_{i}=\ell$ and $c_{i}=c$ for all $i\in[n]$. Note that each item $i\in[n]$ still has objective function values $v_{i}$ and $f_{i}$. This special case is denoted as pu. pu can be solved in polynomial time (Hellion et al., 2012). However, no tight formulation of the problem is known. Hence the proposed perfect compact formulation contributes in this direction.
Constantino Constantino (1998) shows that solving pu, where $\ell_{i}$ can vary with $i\in[n]$, is in general $\mathsf{NP}$-hard. This paper also analyzes a setting identical to pu and obtains two families of valid inequalities. Using both families, it is possible to describe the convex hull when $c\geq d$. Agra and Constantino (2006) study the case where the $y$ variables are general integers and unbounded.
Dash et al. Dash et al. (2016) study the convex hull of the continuous knapsack set. Among other results, they prove that the convex hull with two integer and one bounded continuous variables is described using inequalities for a knapsack set with two integer variables and the ones for a continuous knapsack set with two integer and one continuous variables that is unbounded. We use a special case of this result below.
@@
By Theorem 4.4. in Dash et al. (2016), we know that $\operatorname{conv}\{(\alpha,\gamma)\in\mathbb{R}_{\geq}\times\mathbb{Z}_{\geq}\mathrel{\mathop{\ordinarycolon}}\alpha+\gamma\geq\delta,\alpha\leq\sigma\}$ is the intersection of three sets, namely, $\operatorname{conv}\{(\alpha,\gamma)\in\mathbb{R}_{\geq}\times\mathbb{Z}_{\geq}\mathrel{\mathop{\ordinarycolon}}\alpha+\gamma\geq\delta\}$, $\{(\alpha,\gamma)\in\mathbb{R}\times\mathbb{R}\mathrel{\mathop{\ordinarycolon}}\alpha\leq\sigma\}$ and $\operatorname{conv}\{(\alpha,\gamma)\in\mathbb{R}\times\mathbb{Z}_{\geq}\mathrel{\mathop{\ordinarycolon}}\gamma\geq\delta-\sigma\}$. In the nontrivial case, the mixed integer rounding inequality $\alpha\geq(\delta-\lfloor\delta\rfloor)(\lceil\delta\rceil-\gamma)$ is needed to describe the first set, and the integer rounding inequality $\gamma\geq\lceil\delta-\sigma\rceil$ is needed to describe the third set (see, e.g., Wolsey, 2020). Adding an upper bound of $\nu$ on the integer variable $\gamma$ does not create fractional extreme points.
Now we know that in every extreme point of the set defined by the trivial inequalities and the valid inequalities (28) and (27), $\sum_{i=1}^{\nu}\psi_{i}$ is integer. If $\psi$ is fractional, then it has at least two entries that are fractional and such a point cannot be an extreme point.
@@
$x_{g}^{gb}+\sum_{i=g+1}^{n}\left(d-(b+1)\ell-\left\lfloor\frac{(d-(b+1)\ell}{c}\right\rfloor c\right)y^{gb}_{i}\geq$
$\left(\ell+\left\lceil\frac{d-(b+1)\ell}{c}\right\rceil\times\left(d-(b+1)\ell-\left\lfloor\frac{(d-(b+1)\ell}{c}\right\rfloor c\right)\right)z^{gb}$ $\forall g\in\mathcal{G},b\in B^{g},$
@@
$X^{g1}(b)=\left\{(x,y)\in\mathbb{R}^{n}\times\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}\sum_{i=1}^{g-1}y_{i}=b,y_{g}=1\right\}$
@@
after setting $\nu=n-g$, $\delta=\frac{d-(b+1)\ell}{c}$, $\sigma=\frac{c-\ell}{c}$ and $\alpha=\frac{x_{g}-\ell}{c}$ in Lemma 6.1. Using the result on the convex hull of the union of polyhedra (Conforti and Wolsey, 2008; Balas, 1998), we obtain the tight formulation.